# Attention-based Shape-Deformation Networks for Artifact-Free Geometry Reconstruction of Lumbar Spine from MR Images

### v1

@@ Abstract
Lumbar disc degeneration, a progressive structural wear and tear of lumbar intervertebral disc, is regarded as an essential role on low back pain, a significant global health concern. Automated lumbar spine geometry reconstruction from MR images will enable fast measurement of medical parameters to evaluate the lumbar status, in order to determine a suitable treatment. Existing image segmentation-based techniques often generate erroneous segments or unstructured point clouds, unsuitable for medical parameter measurement. In this work, we present TransDeformer: a novel attention-based deep learning approach that reconstructs the contours of the lumbar spine with high spatial accuracy and mesh correspondence across patients, and we also present a variant of TransDeformer for error estimation. Specially, we devise new attention modules with a new attention formula, which integrates image features and tokenized contour features to predict the displacements of the points on a shape template without the need for image segmentation. The deformed template reveals the lumbar spine geometry in the input image. We develop a multi-stage training strategy to enhance model robustness with respect to template initialization. Experiment results show that our TransDeformer generates artifact-free geometry outputs, and its variant predicts the error of a reconstructed geometry. Our code is available at https://github.com/linchenq/TransDeformer-Mesh.
@@ 1 Introduction
Geometry reconstruction of the anatomical structures from medical images helps to improve clinical outcomes such as surgical planning accuracy, treatment efficacy, and enhanced disease diagnosis and monitoring [1, 2]. Low back pain is a prevalent global health issue associated with activity limitation and absenteeism from work [3, 4], and lumbar disc degeneration, which is the gradual deterioration of the lumbar intervertebral disc, plays a significant role in the onset of this issue [5]. Magnetic Resonance Imaging (MRI) is instrumental in identifying morphological changes and revealing the internal structure of tissues, which is recognized as a key method for investigating disc degeneration [6]. Consequently, it is essential to model lumbar spine morphology and conduct a quantitative analysis of degeneration cascade using a precise geometric representation [7] that is reconstructed from lumbar spine MR images. Due to individual variations, manual geometry annotation is laborious, and therefore automated image analysis methods are desired for clinical applications.
Figure 1: Comparision of our geometry reconstruction approach (right) and the existing image segmentation approaches (left) for lumbar spine MR image analysis.
@@
In contrast to segmentation masks, a mesh (i.e., a structured contour) of the lumbar spine is a more precise representation that depicts the anatomy with nodes/points and elements (i.e., point connectives), as shown in Fig. 1. This mesh representation facilitates computation of the medical parameters that reflect disc degeneration status [7, 9]. Moreover, if meshes with the same topology are used for all patients, mesh correspondence across patients will be established, thus supporting consistent geometric analysis using the same reference anatomical structure [10]. Despite the advantages over segmentation masks, automated geometry reconstruction of the lumbar spine with a mesh representation remains a challenge because of anatomical complexity and variations.
In this study, we propose two new deep neural networks, named UNet-DeformSA and TransDeformer, to reconstruct lumbar spine geometries from 2D sagittal MR images by deforming a mesh template, therefore enabling consistent definition and measurement of the medical parameters related to lumbar disc degeneration. By using new attention modules with a new atten formula, the geometry outputs of our networks are artifact-free. In addition, we propose a third network to estimate the errors of the geometries reconstructed by a network, which facilitates quality control for clinical use.
Due to the 10-page limit, we keep the following sections to be as concise as possible.
@@ 2 Related Works
In our application, the geometry of a lumbar spine is represented by a 2D mesh (see Fig. 1,), i.e., a structured contour of the 11 lumbar spine objects including six vertebrae and five intervertebral discs. Throughout this paper, lumbar spine geometry, shape, and contour are used interchangeably.
As briefly mentioned in the introduction section, in ideal scenarios, the object masks from an image segmentation model could serve as an intermediate representation, from which a mesh could be extracted. In a recent study [11], we evaluated fifteen image segmentation models for our application, (e.g, UNet++ [12], TransUnet [13], Swin-Unet [14], BianqueNet[15]). Although some of these image segmentation models perform well with Dice scores $>$ 0.9 on average, segmentation artifacts are often present in the segmentation masks, which may be (1) extra areas not belonging to any discs or vertebrae, (2) assigning the same class label to two different discs, and (3) broken area of a disc or vertebrae. A mesh extracted from the segmentation masks with artifacts has the same artifacts and therefore is useless. Thus, the segmentation-and-meshing approach do not work for our application.
A template deformation based approach for geometry reconstruction is a better choice for our application, because it ensures mesh correspondence across patients and therefore enables consistent definitions of medical parameters on a reference anatomical structure (i.e., the template). This type of approaches [16] can be traced back to a method known as active appearance model (AAM) [17] that deforms a shape template with the guidance of the statistical model of object geometry and appearance. It is known that the AAM method will not work if images have complex textures and weak edges. To handle complex data, in the pre-deep learning era, classic machine learning (ML) methods were used for template initialization and nonrigid deformation. For example, in the application of heart geometry modeling from CT images [18], a template was initialized by an object detector and then deformed by using an object boundary detector [18], where the detectors were built on probabilistic boosting-tree techniques. In the application of aortic valve leaflet geometry modeling from CT images [19], a template was deformed by using linear coding with a dictionary of representative shapes. Deep learning (DL) has significantly changed the landscape of medical image analysis [2]. In the next few paragraphs, we provide a brief review of the existing medical DL approaches for geometry reconstruction with template deformation.
@@
Instead of treating a geometry as a graph, some studies regards geometry reconstruction as a template-image registration task, wherein a diffeomorphic displacement field [23] is generated by a network to transform an initial template mesh to the target. For example, Daniel et al. proposed DeepCarve that predicted a regular-grid displacement field by a UNet and used it to deform a pre-defined template of left-ventricle and aorta complex on a dataset of 80 cardiac CT scans [24]. As will be shown in our experiments, such a UNet-Disp approach is very sensitive to template initialization, such that some perturbation on initial position of the template may aggravate the quality of the output geometry.
Recently, Transformers utilizing an attention mechanism have been adopted for image classification and segmentation [25] The output of a scaled dot-product attention in Transformers is mathematically formulated as:
@@
Linear projections are applied to the input vectors, resulting in query ($Q$), key ($K$), and value ($V$) matrices. $d_{k}$ is the dimension of a query or key. Unlike convolution which is akin to uniform filtering across different locations on the input, the attention mechanism in Transformers operates as an adaptive filter where its weights are determined by interrelations between each paired tokens.
@@ 3 Methods
As illustrated by Fig. 1, the goal of our application is to reconstruct lumbar spine geometries from 2D sagittal MR images and then measure the medical parameters that are often used for lumbar disc degeneration assessment. For this purpose, the geometry of a lumbar spine is represented by a 2D mesh, i.e., a structured contour of the 11 lumbar spine objects including 6 vertebrae and 5 intervertebral discs.
We designed two novel networks for lumbar spine geometry reconstruction, named UNet-DeformSA and TransDeformer. The UNet-DeformSA model has a UNet backbone and a shape self-attention (SSA) mechanism that sets it apart from the existing models. The TransDeformer model gets rid of the UNet backbone by using cross-attention between shape and image as well as shape self-attention and image self-attention, which further improves the performance. The attention mechanisms in ViT and its successors [26] assume that tokens are located on a regular grid and therefor could not be used in our networks. To enable the new attention modules in our networks, we developed new equations for attention with relative position embedding. Furthermore, we modified TransDeformer for error estimation, i.e., to estimate the error of a reconstructed geometry. In the following sections, we provide the details of our models. For information not revealed due to the 10-page limit, we refer the reader to the source code.
@@ 3.1 Attention with Relative Position Embedding
In the above equation, $W_{Q}$, $W_{K}$, $W_{V}$ are individual linear projections to generate the query, key and value embeddings. We note that in the implementation, the linear projections can be replaced by MLPs. $R_{1}$ and $R_{2}$ are two matrices encoding relative positions between tokens. $d$ is the embedding dimension of a single attention head.
To better explain Eq.(2), we describe this attention from the perspective of individual tokens. We incorporate the sinusoidal functions to capture the relative position information. which is illustrated as:
@@ 3.3 TransDeformer
The architecture of the TransDeformer model is depicted in Fig. 3. Instead of using a UNet, TransDeformer has two groups of CNN layers serving as two feature extractors at two spatial resolutions (512$\times$512 and 128$\times$128). Given the input template shape, the image sampling module will extract image features at each point of the template shape from the feature maps at the lower resolution. Then, image self-attention, shape self-attention, and shape-to-image attention will be performed to cross-examine and fuse information gathered from individual image patches of the feature maps and points of the template. The final output of the shape self-attention module is used to predict the displacement vectors at the points of the template shape, and then the template is deformed to obtain an “intermediate” shape that is further refined by using the higher resolution feature maps, and this refinement is done through a combination of an image sampling layer and a shape self-attention module that outputs adjustments to the “intermediate” displacement vectors. The refined displacement vectors produce the final output shape. The major components of the TransDeformer model are explained with more details in the following sub-sections.
Figure 3: Structure of TransDeformer.
@@ 3.3.1 Image Self-Attention (ISA) Module
The architecture is depicted in Fig. 4. The ISA module aims to produce the sequential image features with the aggregation of global context from the input CNN feature maps. The CNN feature maps are divided into non-overlapping patches [27], thereby enabling contextual dependencies across patches as opposed to individual pixels. A patch is a token in the attention mechanism described in Section A. A mesh grid, $G\in{\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times 2}}$ is used for position embedding, where $P$ is the number of patches, $H$ and $W$ are spatial height and width of a patch. Each patch is centered on one of the mesh grid points. The mesh grid coordinates are normalized between -1 and 1. The ISA module has multiple self-attention layers to generate a sequence of feature maps, enabling conceptual representations with increasingly long-range associate information. Each of the feature maps is converted to patches/tokens in a matrix ${Y_{I}}\in{\mathbb{R}^{L\times E}}$, where $L$ is the number of patches/tokens and $E$ is the embedding/feature dimension.
@@ 3.4 Loss Function, Training and Inference Strategies
Let $\overline{M_{i}}$ and $M_{i}$ denote the predicted and the ground truth binary segmentation masks of the $i$-th class, respectively, where $i\in{\left\{0,1,...,n\right\}}$ and class-0 is the background. $w_{i}$ is a nonnegative scalar inversely proportional to the area of the $i$-th class, with the constraint $\sum_{i=1}^{n}w_{i}=1$. The segmentation quality term $\mathcal{L}_{\text{seg}}$ is formed as:
@@
We proposed a three-stage training strategy to improve accuracy and robustness to template initialization.
@@ 3.4.3 Training Stage 3
Our models have intermediate shape outputs: UNet-DeformsA has three intermediate shape outputs, and TransDeformer has one intermediate shape output. Our loss function consider not only the final output but also the intermediate shape outputs. The loss of geometry quality at each stage is a combination of the loss terms of the intermediate and final shape outputs.
@@ 3.5 Error Estimation Using a Modified TransDeformer
A model for error estimation takes an image and a shape $\hat{S}$ as the inputs, and predicts the difference between the input shape $\hat{S}$ and the ground-truth shape $S$ of the object in the image. If the shape input to the error estimation model is the output from a shape-reconstruction model (e.g., UNet-DeformSA), then the error estimation model provides an estimation of the accuracy of the predicted shape from the shape-reconstruction model.
@@ 4.1 Dataset
Our dataset [11] consists of de-identified lumbar spine MR images of 100 patients. Three experts annotated and reviewed each mid-sagittal MR image to ensure accuracy and consistency. Each mid-sagittal MR image was manually marked with boundaries and landmarks of the lumbar vertebrae and discs, which is guided by the established protocol [7]. The images are resized to 512 $\times$ 512. Each image is accompanied with 12 segmentation masks, corresponding to the background, 6 vertebrae (named L1, L2, L3, L3, L4, L5, and S1), and 5 discs (named D1, D2, D3, D4, and D5). We split the entire dataset into 70 training samples, 10 validation samples, and 20 test samples. A special data augmentation method was applied to the original dataset, resulting in 7000 training samples, 250 validation samples, and 2500 test samples. The data augmentation method is described in details in our technical report [11]. Our experiments were conducted on the augmented datasets that are publicly available [11].
@@ 4.2.1 UNet-GCN Model
This model was originally proposed by Fanwei et al. [22] for reconstructing the geometries of human heart from 3D volumetric images. The model has a UNet backbone pretrained for binary segmentation, and the encoder of the UNet is used as the image feature extractor at different scales. Three GCN modules are used to deform a template sequentially by using image features sampled at the points of the template. For our application, we followed the original design as close as possible, and made necessary adjustment by changing the 3D UNet to a 2D UNet to handle 2D sagittal images of lumbar spline. The UNet-GCN model uses GCN layers to fuse information from neighbour points for displacement prediction. Our UNet-DeformSA model has shape self-attention modules to integrate information from different points that could be far away from each other (i.e., utilizing long-range dependency), which is the key difference from and advantage over the UNet-GCN model.

### v2

@@ Abstract
Lumbar disc degeneration, a progressive structural wear and tear of lumbar intervertebral disc, is regarded as an essential role on low back pain, a significant global health concern. Automated lumbar spine geometry reconstruction from MR images will enable fast measurement of medical parameters to evaluate the lumbar status, in order to determine a suitable treatment. Existing image segmentation-based techniques often generate erroneous segments or unstructured point clouds, unsuitable for medical parameter measurement. In this work, we present TransDeformer: a novel attention-based deep learning approach that reconstructs the geometry of the lumbar spine with high spatial accuracy and mesh correspondence across patients, and we also present a variant of TransDeformer for error estimation. Specially, we devise new attention modules with a new attention formula, which integrate image features and tokenized contour features to predict the displacements of the points on a shape template without the need for image segmentation. The deformed template reveals the lumbar spine geometry in an image. Experiment results show that our TransDeformer generates artifact-free geometry outputs, and its variant predicts the error of a reconstructed geometry. Our code is available at https://github.com/linchenq/TransDeformer-Mesh.
@@ 1 Introduction
Geometry reconstruction of the anatomical structures from medical images helps to improve clinical outcomes such as disease diagnosis accuracy, surgical planning accuracy, and treatment efficacy [1, 2]. Low back pain is a prevalent global health issue associated with activity limitation and absenteeism from work [3, 4], and lumbar disc degeneration, which is the gradual deterioration of the lumbar intervertebral disc, plays a significant role in the onset of this issue [5]. Magnetic Resonance Imaging (MRI) is instrumental in identifying morphological changes and revealing the internal structure of tissues, which is recognized as a key method for investigating disc degeneration [6]. Consequently, it is essential to model lumbar spine morphology and conduct a quantitative analysis of degeneration cascade using a precise geometric representation [7] that is reconstructed from lumbar spine MR images. Due to individual variations, manual geometry annotation is laborious, and therefore automated image analysis methods are desired for clinical applications.
Figure 1: Comparison of our geometry reconstruction approach (right) and the existing image segmentation approaches (left) for lumbar spine MR image analysis.
@@
In contrast to segmentation masks, a mesh (i.e., a structured contour) of the lumbar spine is a more precise representation that depicts the anatomy with nodes/points and elements (i.e., point connectives), as shown in Fig. 1. This mesh representation facilitates the computation of the medical parameters that reflect disc degeneration status [7, 9]. Moreover, if meshes with the same topology are used for all patients, mesh correspondence across patients will be established, thus supporting consistent geometric analysis using the same reference anatomical structure [10]. Despite the advantages over segmentation masks, automated geometry reconstruction of the lumbar spine with a mesh representation remains a challenge because of anatomical complexity and variations.
In this study, we propose two new deep neural networks, named UNet-DeformSA and TransDeformer, to reconstruct lumbar spine geometries from 2D sagittal MR images by deforming a mesh template, thereby enabling consistent definition and measurement of the medical parameters related to lumbar disc degeneration. By using new attention modules with a new attention formula, the geometry outputs of our networks are artifact-free. In addition, we propose a third network to estimate the errors of the geometries reconstructed by a network, which facilitates quality control for clinical use.
Due to the 10-page limit, we keep the following sections to be as concise as possible. We can expand the reference list if the reviewers consider it necessary.
@@ 2 Related Works
In our application, the geometry of a lumbar spine is represented by a 2D mesh (see Fig. 1), i.e., a structured contour of the 11 lumbar spine objects including 6 vertebrae and 5 intervertebral discs. Throughout this paper, lumbar spine geometry, shape, and contour are used interchangeably.
As briefly mentioned in the introduction section, in ideal scenarios, the object masks from an image segmentation model could serve as an intermediate representation, from which a mesh could be extracted. In a recent study [11], we evaluated fifteen image segmentation models for our application, (e.g., UNet++ [12], TransUnet [13], Swin-Unet [14], BianqueNet[15]). Although some of these image segmentation models perform well with Dice scores $>$ 0.9 on average, segmentation artifacts are often present in the segmentation masks, which may be (1) extra areas not belonging to any discs or vertebrae, (2) the same class label assigned to two different discs, and (3) brokens area of a disc or vertebrae. A mesh extracted from the segmentation masks with artifacts has the same artifacts and therefore is useless. Thus, the segmentation-and-meshing approach do not work for our application.
A template deformation based approach for geometry reconstruction is a better choice for our application, because it ensures mesh correspondence across patients and thereby enables consistent definitions of medical parameters on a reference anatomical structure (i.e., the template). This type of approaches [16] can be traced back to a method known as active appearance model (AAM) [17] that deforms a shape template with the guidance of the statistical model of object geometry and appearance. It is known that the AAM method will not work if images have complex textures and weak edges. To handle complex data, in the pre-deep learning era, classic machine learning (ML) methods were used for template initialization and nonrigid deformation. For example, in the application of heart geometry modeling from CT images [18], a template was initialized by an object detector and then deformed by using an object boundary detector [18], where the detectors were built on probabilistic boosting-tree techniques. In the application of aortic valve leaflet geometry modeling from CT images [19], a template was deformed by using linear coding with a dictionary of representative shapes. Deep learning (DL) has significantly changed the landscape of medical image analysis [2]. In the next few paragraphs, we provide a brief review of the existing medical DL approaches for geometry reconstruction with template deformation.
@@
Instead of treating a geometry as a graph, some studies regards geometry reconstruction as a template-image registration task, wherein a diffeomorphic displacement field [23] is generated by a network to transform an initial template mesh to the target. For example, Pak et al. proposed DeepCarve that predicted a regular-grid displacement field by a UNet and used it to deform a pre-defined template of left-ventricle and aorta complex on a dataset of 80 cardiac CT scans [24]. As will be shown in our experiments, such a UNet-Disp approach is very sensitive to template initialization, such that some perturbation on initial position of the template may aggravate the quality of the output geometry.
Recently, Transformers utilizing an attention mechanism, have been adopted for image classification and segmentation [25]. The output of a scaled dot-product attention in Transformers is mathematically formulated as:
@@
Linear projections are applied to the input vectors, resulting in query ($Q$), key ($K$), and value ($V$) matrices. $d_{k}$ is the dimension of a query or key. Unlike convolution which is akin to uniform filtering across different locations on the input, the attention mechanism in Transformers operates as an adaptive filter, where its weights are determined by interrelations between each paired tokens.
@@ 3 Methods
As illustrated by Fig. 1, the goal of our application is to reconstruct lumbar spine geometries from 2D sagittal MR images and then measure the medical parameters that are often used for lumbar disc degeneration assessment. For this purpose, the geometry of a lumbar spine is represented by a 2D mesh, i.e., a structured contour of the 11 lumbar spine components including 6 vertebrae and 5 intervertebral discs.
We designed two novel networks for lumbar spine geometry reconstruction, named UNet-DeformSA and TransDeformer. The UNet-DeformSA model (Fig. 2) has a UNet backbone and a shape self-attention (SSA) mechanism that sets it apart from the existing models. The TransDeformer model (Fig. 3) gets rid of the UNet backbone by using cross-attention between shape and image as well as shape self-attention and image self-attention, which further improves the performance. The attention mechanisms in ViT and its successors [26] assume that tokens are located on a regular grid and therefor could not be used in our networks. To enable the new attention modules in our networks, we developed new equations for attention with relative position embedding. Furthermore, we modified TransDeformer for error estimation, i.e., to estimate the error of a reconstructed geometry. In the following sections, we provide the details of our models. For information not revealed due to the 10-page limit, we refer the reader to the source code.
@@ 3.1 Attention with Relative Position Embedding
$W_{Q}$, $W_{K}$, $W_{V}$ are individual linear projections to generate the query, key and value embeddings. We note that in the implementation, the linear projections can be replaced by MLPs. $R_{1}$ and $R_{2}$ are two matrices encoding relative positions between tokens. $d$ is the embedding dimension of a single attention head.
To better explain Eq.(2), we describe this attention from the perspective of individual tokens. We incorporate the sinusoidal functions to capture the relative position information, which is illustrated as:
@@ 3.3 TransDeformer
The architecture of the TransDeformer model is depicted in Fig. 3. Instead of using a UNet, TransDeformer has two groups of CNN layers serving as two feature extractors at two spatial resolutions (512$\times$512 and 128$\times$128). Given the input template shape, the image sampling module extracts image features at each point of the template shape from the feature maps at the lower resolution. Then, image self-attention, shape self-attention, and shape-to-image attention will be performed to cross-examine and fuse information gathered from individual image patches of the feature maps and points of the template. The final output of the shape self-attention module is used to predict the displacement vectors at the points of the template shape, and then the template is deformed to obtain an “intermediate” shape that is further refined by using the higher resolution feature maps, and this refinement is done through a combination of an image sampling layer and a shape self-attention module that outputs adjustments to the “intermediate” displacement vectors. The refined displacement vectors produce the final output shape. The major components of the TransDeformer model are explained with more details in the following sub-sections.
@@ 3.3.1 Image Self-Attention (ISA) Module
The architecture is depicted in Fig. 4. The ISA module aims to produce the sequential image features with the aggregation of global context from the input CNN feature maps. The CNN feature maps are divided into non-overlapping patches [27], thereby enabling contextual dependencies across patches as opposed to individual pixels. A patch is a token in the attention mechanism described in Section A. A mesh grid, $G\in{\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times 2}}$ is used for position embedding, where $P$ is the number of patches, $H$ and $W$ are spatial height and width of a patch. Each patch is centered on one of the mesh grid points. The mesh grid coordinates are normalized between -1 and 1. The ISA module has multiple self-attention layers to generate a sequence of feature matrices, enabling conceptual representations with increasingly long-range associate information. Each row of a feature matrix ${Y_{I}}\in{\mathbb{R}^{L\times E}}$ is a token, where $L$ is the number of patches/tokens and $E$ is the embedding/feature dimension.
@@ 3.4 Loss Function, Training and Inference Strategies
Let $\overline{M_{i}}$ and $M_{i}$ denote the predicted and the ground truth binary segmentation masks of the $i$-th class respectively, where $i\in{\left\{0,1,...,n\right\}}$ and class-0 is the background. $w_{i}$ is a nonnegative scalar inversely proportional to the area of the $i$-th class, with the constraint $\sum_{i=1}^{n}w_{i}=1$. The segmentation quality term $\mathcal{L}_{\text{seg}}$ is formed as:
@@
Also, we proposed a three-stage training strategy to improve spatial accuracy on meshes and robustness to template initialization.
@@ 3.4.3 Training Stage 3
Our models have intermediate shape outputs: UNet-DeformsA has three intermediate shape outputs, and TransDeformer has one intermediate shape output. Our loss function considers not only the final output but also the intermediate shape outputs. The loss of geometry quality at each stage is a combination of the loss terms of the intermediate and final shape outputs.
@@ 3.5 Error Estimation Using a Modified TransDeformer
A model for error estimation takes an image and a shape $\hat{S}$ as the inputs, and predicts the difference between the input shape $\hat{S}$ and the ground-truth shape $S$ of the object in the image. The shape input to the error estimation model could be the output from a shape-reconstruction model (e.g., UNet-DeformSA).
@@ 4.1 Dataset
Our dataset [11] consists of de-identified lumbar spine MR images of 100 patients. Three experts annotated and reviewed each mid-sagittal MR image to ensure accuracy and consistency. Each mid-sagittal MR image was manually marked with boundaries and landmarks of the lumbar vertebrae and discs, which is guided by the established protocol [7]. The images are resized to 512 $\times$ 512. Each image is accompanied with 12 segmentation masks, corresponding to the background, 6 vertebrae (named L1, L2, L3, L3, L4, L5, and S1), and 5 discs (named D1, D2, D3, D4, and D5). We split the entire dataset into 70 training samples, 10 validation samples, and 20 test samples. A special data augmentation method was applied to the original dataset, resulting in 7000 training samples, 250 validation samples, and 2500 test samples. The data augmentation method is described in details in our technical report [11]. Our experiments were conducted on the augmented datasets that are publicly available [11]. Data diversity is illustrated in Fig. 8 by comparing the template with the ground-truth meshes, for which the centroids are aligned.
Figure 8: Distribution of Averaged Point-to-Point Distance (APPD) between the template mesh and the ground truth meshes of the patients in the test set. APPD is defined in Section 4.4.
@@ 4.2.1 UNet-GCN Model
This model was originally proposed by Kong et al. [22] for reconstructing human heart geometries from 3D volumetric images. The model has a UNet backbone pretrained for binary segmentation, and the encoder of the UNet is used as the image feature extractor at different scales. Three GCN modules are used to deform a template sequentially by using image features sampled at the points of the template. For our application, we followed the original design as close as possible, and made necessary adjustment by changing the 3D UNet to a 2D UNet to handle 2D sagittal images of lumbar spline. The UNet-GCN model uses GCN layers to fuse information from neighbour points for displacement prediction. Our UNet-DeformSA model has shape self-attention modules to integrate information from different points that could be far away from each other (i.e., utilizing long-range dependency), which is the key difference from and advantage over the UNet-GCN model.